### v1

# Conceptualizing Uncertainty

Abstract
Uncertainty in machine learning refers to the degree of confidence or lack thereof in a model’s predictions. While uncertainty quantification methods exist, explanations of uncertainty, especially in high-dimensional settings, remain an open challenge. Existing work focuses on feature attribution approaches which are restricted to local explanations. Understanding uncertainty, its origins, and characteristics on a global scale is crucial for enhancing interpretability and trust in a model’s predictions. In this work, we propose to explain the uncertainty in high-dimensional data classification settings by means of concept activation vectors which give rise to local and global explanations of uncertainty. We demonstrate the utility of the generated explanations by leveraging them to refine and improve our model.  [^1] Code will be freely available here (upon acceptance): https://github.com/robertsi20/Conceptualizing-Uncertainty.
Keywords:
Explainable UncertaintyConcept-based ExplanationsXAI.
1 Introduction
While advances in deep learning in the last years have led to impressive performance in many domains, such models are not always reliable, particularly when it comes to generalizing to new environments or adversarial attacks. To improve on that, numerous methods have been developed in the field of explainable artificial intelligence (xAI) [5] to provide insights into model behavior and facilitate actionable modifications. However, the majority of methods focus on explaining model predictions, which can help understand misclassifications but do not explicitly address predictive uncertainty(See Figure 1). Understanding uncertainty is crucial for detecting potential model weaknesses, particularly in dynamic environments.
Since uncertainty quantification is useful in various applications, including active learning [20], classification with rejects [17], adversarial example detection [26], and reinforcement learning [24], a significant body of work aims to improve the quantification of predictive uncertainty using Bayesian deep learning (BDL) and approximations thereof [15, 9, 14]. In contrast, the literature on understanding the sources of uncertainty for a given model via explanations is limited, focusing on methods for feature attribution [28, 27] (see section 2.4 for more related work). While these approaches do provide valuable insights for individual input data points, they are restricted to local explanations and do not provide high-level insights on a global dataset scale, e.g., for image datasets. Also, another limitation of feature attribution methods, known from the prediction attribution setting and prevalent in the image domain, is that these point to which part of the input the model considers as important, but not why [13]. Here, methods aiming to automatically extract concept activation vectors (CAVs) [11, 1] and attributing them back to the input do provide both, information on what the model seems to perceive in a part of the input, and global dataset information, both by inspecting the learned concepts.
In this work, (i) we propose a pipeline to enable concept-based explanations for predictive uncertainty, providing both local and global explanations for sources of uncertainty, and (ii) we demonstrate the correctness and usefulness of the learned concepts as well as the potential to perform actionable interventions based on these in a series of downstream tasks, including the automatic detection of different types of novelty in new environments, improving uncertainty-based rejections and detecting gender bias in language models.
We proceed by describing foundations for our contributions, including related work on uncertainty explanations. This is followed by our proposed pipeline in section 3, experiments in 4, a discussion on limitations in 5 and a conclusion.
Figure 1: A White Wolf classified as Samoyed by an ImageWoof-adapted ResNet50 classifier. We generated the concept attribution map explaining the prediction (left) versus explaining the uncertainty (right) with our proposed pipeline. Correspondingly, we display three important concepts for this image using each explanation approach. The prediction explanation suggests the neural net arrived at its decision due to the detected concepts of a Samoyed’s ears, face, and legs, while the uncertainty explanation offers that the model is uncertain about its prediction because of the detected concept of a wolf’s head, fur, and legs.
2 Background
In this section, we list the relevant fundamentals, including our problem formulation, background on uncertainty quantification and concept activation vectors, as well as related work on explaining uncertainty.
2.1 Setup and Problem Formulation
Given a trained classification model $\mathbf{M}$, e.g., a deep convolutional network, and a dataset of $n$ data points $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{n}\}$, usually not seen during training. In contrast to a vast amount of literature focusing on explaining predictions of such models [5], we aim for explaining their predictive uncertainty, thereby aiming to understand its sources. Our work differs from existing recent research on that topic, e.g., [5], in that we aim for global explanations.
For the following Bayesian formalization, we also require the training data set $\mathcal{D}$ of $\mathbf{M}$, which, however, is not used by our method.
2.2 Quantifying Uncertainty
Predictive uncertainty is typically defined over the predictive distribution $p(y|\mathbf{x},\mathcal{D})$ [15, 9], which in our case is over possible labels $y$. In a Bayesian setting, where the parameters $\boldsymbol{\theta}$ of our classification model $\mathbf{M}$ are random variables, $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]$ requires computing the expectation $\mathbb{E}$ over the posterior $p(\boldsymbol{\theta}|\mathcal{D})$, which in practice is usually intractable. Accordingly, various methods for posterior approximation have been introduced such as ensemble methods [21], Monte Carlo (MC) dropout [14] and Laplace Approximations [6].
Different measures for predictive uncertainty are defined in literature [15, 9]. We summarize the ones we are using in our experiments in the following.
Total Uncertainty based on Shannon Entropy:
$u_{t}(\mathbf{x})\coloneq\mathcal{H}[p(y|\mathbf{x},\mathcal{D})]=-\sum_{y\in Y}p(y|\mathbf{x},\mathcal{D})\log_{2}p(y|\mathbf{x},\mathcal{D})$ (1)
Aleatoric and Epistemic Uncertainty based on the decomposition of $u_{t}$:
$u_{a}(\mathbf{x})\coloneq\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}\big{[}\mathcal{H}[p(y|\mathbf{x},\boldsymbol{\theta})]\big{]},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u_{e}(\mathbf{x})\coloneq u_{t}(\mathbf{x})-u_{a}(\mathbf{x})$ (2)
In order to approximate the above measures, we utilize MC dropout to collect a set predictions $\{p(y|\mathbf{x},\boldsymbol{\theta}_{i})\}_{i=1}^{N}$ and approximate the posterior predictive $p(y|\mathbf{x},\mathcal{D})=\mathbb{E}_{p(\boldsymbol{\theta}|\mathcal{D})}[p(y|\mathbf{x},\boldsymbol{\theta})]\approx\frac{1}{N}\sum_{i}^{N}p(y|\mathbf{x},\boldsymbol{\theta}_{i}).$ We refer to $\hat{u}_{t},\hat{u}_{a},\hat{u}_{e}$ when utilizing this approximation in $u_{t},u_{a},u_{e}$, respectively.
2.3 Concept Activation Vectors
Concept Activation Vectors (CAVs) aim for human interpretability with respect to understanding black-box model predictions [19, 11, 16]. Recent approaches based on Nonnegative Matrix Factorization (NMF) [13, 18] have been shown to demonstrate superior qualitative and quantitative properties of the resulting concepts[11]. They work by embedding patches of inputs into a non-negative activation space of a pre-trained model and applying NMF to decompose the embedded data matrix $\mathbf{A}$ into a product of non-negative matrices $\mathbf{W}$ and $\mathbf{V}$, solved by reconstructing $\mathbf{A}$, i.e.,  $(\mathbf{W},\mathbf{V})=\arg\min_{\mathbf{W}\geq 0,\mathbf{V}\geq 0}\|\mathbf{A}-\mathbf{W}\mathbf{V}^{\top}\|^{2}_{F}$. The decomposition yields: $\mathbf{V}$ the dictionary of concepts (or concept bank) and $\mathbf{W}$ a reduced representation of $\mathbf{A}$ according to the basis $\mathbf{V}$. To attribute importance, they make use of a sensitivity analysis technique known as total Sobol Indices, which captures the effects of a concept along with its interactions on the model’s output by considering the variance fluctuations by perturbing $\mathbf{W}$.
2.4 Related Work on Explaining Uncertainty
While there does exist a plethora of xAI methods for explaining the prediction of classification models including several local and global approaches [5], methods for explaining the source of uncertainty have only developed in the last years. So far, these have focused on local feature attribution explanations, including explaining uncertainty with shapley values [28], with gradient-based methods [27], with counterfactuals [2] and by taking second order effects into account [4]. In contrast, we aim for explanations beyond feature attribution, that provide also global explanations of uncertainty, which enable an overview of characteristics of uncertain predictions. Also, concerning image datasets, rather small ones consisting of MNIST and CIFAR are used in [28, 27, 2]. Only [4] is applied to CelebA, containing larger images. We apply our approach to images of ImageNet.
Another, less related line of work, aims for uncertainty sets, dealing with the problem of uncertainty of explanations [22].
3 Proposed Pipeline
Figure 2: Our proposed pipeline for uncertainty explanation using CAVs.
We propose to explain predictive uncertainty by means of CAVs computed on a local level – for each data point – or aggregated to obtain a global explanation. In Figure 2, we illustrate our proposed pipeline which aims to characterize and explain uncertainty using extracted concepts using high-dimensional data $\mathcal{D}$. Given also a classification model $\mathbf{M}$, and an uncertainty measure $u$, e.g., $u\coloneq\hat{u}_{t}$, we arrange our pipeline into 4 steps:
Step 1 - We use the model $\mathbf{M}$ to classify the inputs and compute $u(\mathbf{x})$ for each data point, by leveraging an approximation technique such as MC Dropout.
Step 2 - To apply concept activations, we need to specify a probabilistic classification task. In our setup, a canonical choice is the grouping induced by uncertain (UNC) and certain (CER) samples. To obtain a reasonable decision boundary, we observe that $u$ will map CER to small numbers and UNC to large numbers. We therefore expect that if applied to all data points, $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$ can be described by a mixture model with two components. For simplicity, we assume a Gaussian Mixture Model (GMM) with two components which we train on $\{u(\mathbf{x}_{i})\}_{i=1}^{n}$. By our assumption we expect the component with the larger mean to correspond to the UNC samples. We thus obtain the classification model $f$ by considering the conditional probability which takes on a sigmoid shape. To counteract potential class imbalance, i.e., an imbalance between CER and UNC in the training data, we set the weights of the mixture component to equal size after training.
Step 3 - To generate the concepts, we embed the data using a foundation model $g$ into an activation space with the condition that for each $\mathbf{x}_{i}$, $g(\mathbf{x}_{i})\geq 0$ (e.g., after a ReLU layer), and then we train one NMF on patches from $\{g(\mathbf{x}_{i})|\mathbf{x}_{i}\in\text{UNC}\}$ and another NMF on patches from $\{g(\mathbf{x}_{i})|\mathbf{x}_{i}\in\text{CER}\}$, producing two concept banks, $\mathbf{V}_{\text{UNC}}$ and $\mathbf{V}_{\text{CER}}$. Thus, we can represent each $g(\mathbf{x}_{i})$ as a linear combination of the concepts in $\mathbf{V}_{\text{\text{UNC}}}$ or $\mathbf{V}_{\text{\text{CER}}}$, with scaling factors $\mathbf{W}_{i}$.
Step 4a - To estimate the importance of the concepts in $\mathbf{V}_{\text{\text{UNC}}}$ and $\mathbf{V}_{\text{\text{CER}}}$, we utilize the Sobol Indices [12, 13, 18], using $f$ as the function of interest.
Step 4b - Repeating Step 4a for every data point, we obtain a local importance score $e_{l}(g(\mathbf{x}_{i}))\in\mathbb{R}^{d}$ with $d$ concepts. Additionally, we supplement the local importances with an attribution map[13, 18] indicating where the important concepts are detected in the input. We further augment the local explanations $e_{l}$ with consistent global explanations. The global importances can be computed as in literature by averaging over $e_{l}(g(\mathbf{x}_{i}))$ for points predicted in UNC and CER, respectively, producing $e_{\text{UNC}}$ and $e_{\text{CER}}$.
4 Experiments
Since uncertainty arises from multiple sources, establishing a ground truth for experimentation can be challenging. Therefore, we aim to demonstrate the validity and usefulness of our uncertainty explanations by integrating them into a classification with reject options setting. Additionally, we illustrate how our concepts capture different sources of uncertainty, aiding human decision-makers in constructing effective re-training sets. Finally, we show that our concepts can reveal potential biases with respect to sensitive groups in a downstream task.
4.1 Rejecting Uncertain Points with Concepts
Figure 3: (left) Accuracy-Rejection ($\uparrow$) and (right) OOD-Rejection Curves ($\downarrow$).
We demonstrate that our proposed explanations enable actionable interpretations by utilizing the learned concepts for rejection decisions in a classification setting. In particular, given a trained model $\mathbf{M}$ and a set of new data points $\mathbf{X}$, we apply our proposed approach to generate concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$.
Now we consider the question: Can we improve a basic rejection strategy, which rejects a fixed amount of points based on predictive uncertainty, by leveraging our explanations? For the latter, we consider two strategies: 1. Concept-only rejection: we identify for each input $\mathbf{x}_{i}$ the most strongly activated concept $c^{*}=\arg\max\mathbf{W}_{i}$ utilizing the combined concept bank $[\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}]$ and determine the global importance of $c^{*}$. We then reject those points first, for which $c^{*}\in e_{\text{UNC}}$ and with highest $e_{\text{UNC}}$ value. This leads to inputs associated with globally uncertain concepts being rejected first, while those linked to globally certain concepts being retained longer. 2. Weighted rejection: We adapt the previous strategy, by weighting the uncertainty output $f(\mathbf{x}_{i})$ with $+1$ or $-1$, depending on whether $c^{*}\in e_{\text{UNC}}$ or $c^{*}\in e_{\text{CER}}$ and again reject according to highest value.
To evaluate the effectiveness of these methods, we use a pre-trained ResNet-50 classifier as the base model. We randomly sample images from 10 out of 20 ImageNet classes (ImageWoof + Imagenette) and also include out-of-distribution (OOD) samples from the NINCO dataset [3]. In each run, we use 1,000 points, with 40% being OOD to compute accuracy-rejection curves [23, 25], which depict the accuracy computed on all not-rejected points on the y-axis for different percentage rejection rates (x-axis). We compare our two concept-based strategies, Concept-only and Weighted rejects, to baselines, which reject according to highest predictive uncertainty, using each of $\hat{u}_{t},\hat{u}_{e},\hat{u}_{a}$. We also compute the AUC for each strategy and average these measures over 20 runs. We set the number of concepts to 55 for the certain group and 35 for the uncertain.
As seen in Figure 3 (left), our "weighted" method performs with the highest AUC score computed across 20 runs. To confirm the statistical significance of this result, we conduct a one-sided Wilcoxon signed rank test against the Total Entropy (which in this case is very similar to the Aleatoric curve), obtaining a p-value $<10^{-6}$, indicating ours performing significantly better. Meanwhile, the concept-only strategy performs better for medium rejection rates and worse for higher, as compared to the baselines.
In Figure 3 (right), we plot the percentage of OOD points among the not-rejected ones. Up to approximately 20% rejection, all curves except the Epistemic and Concept-only one, exhibit similar performance. However, beyond this point, both our strategies reject a greater proportion of OOD samples at each step. This signifies why our explanations improve the rejection: the learned concepts pick up on the OOD data and helps in rejecting them.
4.2 Distinguishing Sources for Uncertainty
Figure 4: (left) Motion Blur experiment results ($\uparrow$) and an example noisy image (right) to better visualize the noise and the corresponding concept extracted.
In this experiment, we assess the quality and usefulness of the obtained explanations by showing that the learned concepts allow for a grouping of different sources of uncertainty, such that humans can better identify these sources and make more informed downstream decisions.
In particular, our setting involves a classifier $\mathbf{M}$ that is applied in the wild and is exposed to two different types of noise. We consider structured noise induced by novel classes not seen during training and unstructured noise, like random distortions (e.g., blurring). We want to evaluate in how far our explanations can help to detect the different noise types and filter the instances with unstructured noise out, without having explicit knowledge about the noise beforehand. This can occur in an active learning setup, where a data exploration task could be to filter out unrecognizable noisy images that provide no meaningful learning signal and finetune the model on the remainder.
More precisely, we consider the setting where new data samples $\mathbf{X}$ are available and our explanation pipeline is applied to provide concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$. We aim to evaluate how well $\mathbf{V}_{\text{UNC}}$ captures different noise types, and how well unstructured noise samples $\mathbf{x}_{i}$ can be filtered out using $e_{l}(\mathbf{x}_{i})$ and $\mathbf{W}_{i}$ of $\mathbf{V}_{\text{UNC}}$ concepts. For this purpose, we assume that we can determine the set $nc$ that correspond to unstructured noise concepts in $\mathbf{V}_{\text{UNC}}$ (as a human could visually inspect them). Then, for each data point $\mathbf{x}_{i}$, we sum the local importance $e_{l,nc}(\mathbf{x}_{i})$ or NMF activations $\mathbf{W}_{i,nc}$ and filter out according to highest values, corresponding to the amount of presence of these concepts. We refer to these strategies as Ours (Importance) when using $e_{l,nc}(\mathbf{x}_{i})$, and to and Ours (NMF) for $\mathbf{W}_{i,nc}$.
[…]
Table 1: We report the average AUC score ()($\uparrow$)( ↑ ) over 20 runs for various types of random noise patterns. Our methods using the local importances or reduced NMF representation perform better than the baselines. We test our methods against Gaussian Blurring (G Blur), Salt and Pepper Noise (S and P), Gaussian Noise (G Noise), Motion Blurring (M Blur), Radial Blurring (R Blur), and Wave Noise(Wave)
Method Total Aleatoric Epistemic Ours (Imp) Ours (NMF)
G Blur $79.6\pm 1.5$ $79.5\pm 1.5$ $82.6\pm 1.2$ $\underline{85.6}\pm 1.3$ $\textbf{89.0}\pm 0.2$
SnP Noise $77.2\pm 1.4$ $77.3\pm 1.4$ $76.7\pm 1.7$ $\underline{85.4}\pm 0.9$ $\textbf{88.9}\pm 0.2$
G Noise $82.3\pm 1.3$ $82.5\pm 1.3$ $70.8\pm 2.7$ $\underline{85.1}\pm 1.2$ $\textbf{88.8}\pm 0.3$
M Blur $77.5\pm 2.0$ $77.4\pm 2.0$ $80.4\pm 1.4$ $\underline{87.0}\pm 0.9$ $\textbf{89.2}\pm 0.2$
R Blur $82.2\pm 1.6$ $82.4\pm 1.6$ $73.1\pm 2.9$ $\underline{85.2}\pm 4.4$ $\textbf{86.2}\pm 4.9$
Wave Noise $80.8\pm 7.4$ $80.1\pm 8.0$ $87.8\pm 2.0$ $\textbf{88.3}\pm 1.5$ $\textbf{88.3}\pm 1.5$
Average $79.9\pm 2.2$ $79.9\pm 2.3$ $78.5\pm 6.3$ $\underline{86.1}\pm 1.3$ $\textbf{88.4}\pm 1.1$
The results in Table 1 indicate that our method outperforms uncertainty-based measures for this task. Thereby, using $e_{l}(g(\mathbf{x}_{i}))$ achieves the second-best performance (underlined), while leveraging $\mathbf{W}_{i}$ yields the best results (bolded), with the highest average AUC score. Inspecting Figure 4, we can see that our method using the NMF coefficients (purple) consistently recommends images that would be beneficial to the re-training set while abstaining from recommending the unusable images until they must be chosen. We indicate the true unusable image percentage by the vertical line on the figure. In the ideal case, a method would not recommend any unusable images until it reaches the vertical line. Our method provides an explainable way to automatically select images which contain meaningful structure for domain adaptation.
4.3 Explaining Uncertainty in Language Models for Fairness
In this experiment, we demonstrate our proposed pipeline can also be applied to the natural language domain and that our explanations can capture sensitive group information, which can be used to correct bias in a model’s predictions.
For this purpose, we fine-tune a BERT [10] model on the Bias in Bios dataset [7] which consists of biographies where the task is to predict their corresponding occupation. Before fine-tuning, we incorporate ReLU on the last embedding layer to ensure non-negativity. We then apply our pipeline using a NMF-based concept extraction technique for the text domain [18]. We inspect the "physician" class, by training an NMF on the inputs predicted as such and compute their importances with respect to each group of uncertain and certain points.
We investigate the most important concept of the uncertain samples, number 6, in the following. In Figure 5, to understand what concept 6 represents, we show two excerpts which activate concept 6. Thereby, the intensity of red marks the strength of the activation of concept 6. We can see that female pronouns appear among the highlighted words along with other nouns like "co-founder" and "sleep medicine specialist". Since we know the gender labels of the dataset, we check the Pearson correlation between concept 6 and the labels. Indeed, it is the most correlated with gender at $R=0.3$. We further verify the relevance of this concept for representing gender by excluding it in the NMF reconstruction and applying the occupation classifier. This changes some of the predictions, most notably a large proportion of professors and chiropractors that were falsely predicted as physicians. Even more interesting, the gender distribution among the samples influenced by concept 6 does not align with the gender distribution of said classes, which could be an indication of gender bias in BERT. We evaluate the change in gender bias by computing the equalized odd score before and after our intervention and report an improvement of 0.0027. At first glance this score does not sound impressive, the intervention on the physician class only affects a limited number of samples and thus has a limited effect on global equalized odds. More precisely, the best possible outcome for an intervention on the physician class (fixing all false positives) would have led to an improvement of 0.0069. This demonstrates that the detected relevant concept in the uncertain group encodes gender information in our example, and that removing it can improve fairness.
Figure 5: Text excerpts where activations of concept 6 are highlighted with red.
5 Limitations and Future Work
While our approach performs well in the tasks outlined above, it is not without limitations. Concept-based explanations provide a human-interpretable means of understanding uncertainty in machine learning settings, particularly when the source of uncertainty is visibly discernible. However, they may fail to capture finer-grained pixel-level nuances of uncertainty. In this study, we maintained a fixed patch size when training the NMF, but varying the patch size could potentially capture more localized properties of uncertainty. Investigating the relationship between patch size and its impact on concept-based explanation quality presents a promising direction for future research.
Additionally, our method for estimating concept importance may not be optimal, as evidenced by the superior performance of using NMF activations directly in our experiments. We suspect this performance gap arises from the variance—or lack thereof—of the uncertainty measure. Specifically, if we perturb a concept within an already highly uncertain input, the uncertainty measure may not exhibit significant variation. While we did not explore alternative ways to refine importance attribution in this study, we do plan to do address it in future iterations of this work.
6 Conclusion
We introduced a novel framework for explaining uncertainty using automatically extracted concept activation vectors. Our proposed framework enables both local and global explanations of uncertainty through the use of importance scores and attribution maps. These explanations demonstrate their utility by being incorporated into rejection strategies, aiding design of useful re-training sets, and helping to detect and mitigate bias. Moreover, while concept-based explanations of model predictions can be useful, using CAVs to capture sources of uncertainty not only offers another complementary view into how a model makes its decisions, but also provides interpretable ways to enhance its performance.
7 Acknowledgements
This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101073307.